How do you solve the answer to 3/5x=9

Found 2 solutions by user_dude2008, Alan3354:
Answer by user_dude2008(1862)

3/5x=9

3x=9*5

3x=45

x=45/3

x=15

Answer by Alan3354(69443)

3/5x = 9
Multiply by 5x
3 = 45x
x = 3/45
x = 1/15